А B 0.2 kg wwwwww 90 mm k = 1.2 N/mm

Ring A is drawn by a spring of normal length (without stretching) of 100 mm and k = 1.2 N / mm. The bracelet standing at point A is released as soon as x = 200 mm. Calculate the initial acceleration
Answer #1

when x = 200 mm, length of the spring in the figure, L = sqrt(90^2 + 200^2)

= 219 mm

extension of the spring, delta_x = 219 - 100

= 119 mm

Force exerted by the spring, F = k*delta_x

= 1.2*119

= 142.8 N

angle made by this force horizontal, theta = tan^-1(90/200)

= 24.2 degrees.

initial acceleration, a = Fnetx/m

= F*cos(24.2)/m

= 142.8*cos(24.2)/0.2

= 651 m/s^2 <<<<<<<<<---------------------Answer